# ADGaussian: Generalizable Gaussian Splatting for Autonomous Driving with Multi-modal Inputs

### v1

@@ Abstract
We present a novel approach, termed ADGaussian, for generalizable street scene reconstruction. The proposed method enables high-quality rendering from single-view input. Unlike prior Gaussian Splatting methods that primarily focus on geometry refinement, we emphasize the importance of joint optimization of image and depth features for accurate Gaussian prediction. To this end, we first incorporate sparse LiDAR depth as an additional input modality, formulating the Gaussian prediction process as a joint learning framework of visual information and geometric clue. Furthermore, we propose a multi-modal feature matching strategy coupled with a multi-scale Gaussian decoding model to enhance the joint refinement of multi-modal features, thereby enabling efficient multi-modal Gaussian learning. Extensive experiments on two large-scale autonomous driving datasets, Waymo and KITTI, demonstrate that our ADGaussian achieves state-of-the-art performance and exhibits superior zero-shot generalization capabilities in novel-view shifting. The project page can be found at maggiesong7.github.io/research/ADGaussian/.
Figure 1: We introduce ADGaussian, a generalizable Gaussian framework for street scene reconstruction. Our approach achieves superior performance in both visual and geometric reconstruction. The bottom row illustrates the results of viewpoint shifting, further demonstrating the robustness of our method under varying viewpoint changes.
1 Introduction
@@
Recently, 3D Gaussian Splatting (3DGS) [14] has garnered significant attention in the fields of 3D scene reconstruction and novel view synthesis due to its real-time rendering speed and high-quality output. One key application is the modeling of street scenes from image sequences, which plays a vital role in areas such as autonomous driving.
@@
When modeling urban scenes, some methods follow per-scene optimization techniques [48, 4, 17], notably StreetGaussians [38] that represents dynamic urban street as a set of point clouds equipped with semantic logits and 3D Gaussians. While per-scene optimization approach excels in high-quality reconstruction, it often struggles with expensive training cost and large-range novel view synthesis.
@@
To achieve generalizable street scene reconstruction, most existing methods build upon the architectures of PixelSplat [3] or MVSplat [6]. For instance, GGRt [19] introduces a pose-free architecture to iteratively update multi-view depth map and subsequently estimates Gaussian primitives based on PixelSplat. Similarly, GGS [9] enhances the depth estimations of MVSplat by integrating a multi-view depth refinement module. Nevertheless, multi-view feature matching-based depth estimation may fail in challenging conditions such as texture-less areas, and reflective surfaces. To tackle this issue, the concurrent work DepthSplat [36] combines pre-trained depth features from Depth Anything V2 [40] with multi-view depth estimations for accurate depth regression, where the estimated depth features are further used for Gaussian prediction.
@@
Given the great generalization capability of Depth Anything V2, it is reasonable to extend DepthSplat to urban street scenarios. However, DepthSplat faces specific limitations when applied to these environments. First, the visual rendering quality is constrained by the effectiveness of pre-trained depth models. Additionally, even in cases where depth quality is high, straightforwardly concatenating image and depth features for Gaussian prediction leads to unsatisfactory visual reconstruction in complex autonomous driving situations (see Figure 1).
@@
To this end, we propose a novel multi-modal representation framework aimed at enhancing both geometry modeling and visual rendering in street scenes. This framework leverages two complementary modalities: image visual cues and depth geometry clues. The key insight is to facilitate effective information sharing and joint optimization between different modalities. Specifically, our framework begins by integrating sparse LiDAR depth information as an additional input modality, which provides precise metric scale priors for enhanced geometry reconstruction. Given the image data and sparse depth map, we introduce an effective multi-modal feature matching strategy achieved by a Siamese-style encoder paired with an information cross-attention decoder. This design ensures a cohesive fusion of geometric and appearance information, resulting in well-aligned multi-modal tokens. Subsequently, we employ a multi-scale gaussian decoding model to aggregate multi-scale depth information into both image and depth tokens for the final Gaussian predictions. Notably, our model exhibits superior performance even under large viewpoint changes, as shown in the bottom row of Figure 1. This capability demonstrates the effectiveness of jointly optimizing multi-modal features, which enables superior geometry and texture reconstruction.
@@
  • We present ADGaussian, the first generalizable framework that models street scenes using multi-modal data, specifically color images and depth maps.
@@
  • We perform extensive comparisons on two large-scale driving datasets, showcasing our approach’s state-of-the-art performance and demonstrating the effectiveness of the proposed components.
2 Related Work
2.1 Generalizable 3D Gaussian Splatting
@@
Generalizable Gaussian Splatting [32, 43, 37, 34, 5, 11] aims to learn powerful priors that enable effective generalization across unseen scenes. Existing methods can be broadly categorized into two groups. The first group, including approaches like PixelSplat [3], MVSplat [6], MVSGaussian [20], and SplatterImage [26], predicts per-pixel 3D Gaussian primitives using known camera parameters. The second group of methods [13, 28, 23, 41] proposes to jointly predict camera parameters and 3D representations, eliminating the need for known camera poses. For instance, GGRt [19] employs an Iterative Pose Optimization Network to estimate and iteratively update the relative pose between target and reference images. In street scene modeling, however, camera poses provide critical constraints for determining scene scale and enhancing reconstruction accuracy from video sequences. Moreover, camera poses are readily accessible in street scenes, making them a practical and reliable data resource. Therefore, we choose to leverage posed images for our approach.
Figure 2: Overview of ADGaussian. Given monocular posed image with sparse depth as input, we first extract well-fused multi-modal features through Multi-modal Feature Matching, which contains a siamese-style encoder and a cross-attention decoder enhanced by Depth-guided positional embedding (DPE). Subsequently, the Gaussian Head and Geometry Head, augmented with Multi-scale Gaussian Decoding, are utilized to predict different Gaussian parameters.
2.2 Depth and Gaussian Splatting
@@
Depth quality has been demonstrated to play a pivotal role in Gaussian Splatting. To ensure accurate geometry reconstruction, methods such as GPS-Gaussian [47] and DN-Splatter [29] incorporate additional depth supervision into the optimization process. However, as dense ground truth depth data is often unavailable, an alternative research direction has emerged, leveraging pre-trained depth foundation models [21, 40] to provide reliable geometric cues. For example, Chung et al. [7] rescale pre-trained depth maps using sparse COLMAP points, thereby offering accurate depth constraints for model optimization. Additionally, Flash3D [25] utilizes a frozen off-the-shelf network to estimate the metric depth, which subsequently serves as a direct input to the reconstruction model. To mitigate error propagation, DepthSplat [36] fuses pre-trained depth features with multi-view cost volume features while keeping the pre-trained network learnable. In contrast to previous approaches that focus primarily on geometry improvement, we argue that joint optimization of image and depth features is more critical for achieving high-quality reconstruction.
2.3 LiDAR-Integrated Gaussian Splatting
@@
The integration of LiDAR data has emerged as a widely adopted approach in street scene reconstruction, mainly due to its effectiveness in facilitating geometry learning. The conventional methodologies typically involve two main steps: initializing Gaussians from LiDAR point clouds [15], and further supervising predicted Gaussian positions using LiDAR priors [49, 4, 12, 10]. However, reconstruction quality often degrades in areas where LiDAR measurements are unavailable. To address this limitation, HGS-Mapping [35] extracts matching feature points from adjacent RGB frames and utilizes the estimated spatial positions to initialize areas beyond LiDAR’s coverage. In the same manner, TCLC-GS [46] constructs a hybrid 3D representation by combining LiDAR geometries with image colors, enabling simultaneous initialization of both geometric and appearance attributes of 3D Gaussians. Rather than directly using LiDAR point clouds, we propose leveraging sparse LiDAR depth to bridge the gap between LiDAR and camera data. Furthermore, we integrate depth priors into the training process, achieving joint optimization of depth geometries and image photometric attributes, as opposed to the common practice of initialization alone.
3 Methods
@@
Depth foundation models [1, 2, 39, 40, 42] have been integrated into the Gaussian Splatting to improve geometry reconstruction. However, such framework often suffer from suboptimal rendering quality due to the insufficient interactions between photometric and geometric clues. To address this, we propose ADGaussian, a synchronized multi-modal optimization architecture that combines sparse depth data with monocular images for enhanced street scene modeling.
3.1 Preliminary: Depth Foundation Model for Gaussian Splatting
Recently, works like DepthSplat have investigated the advantages of using a pre-trained depth foundation model for image-conditioned 3D Gaussian reconstruction, capitalizing on its remarkable performance across a variety of real-world datasets. All these methods utilize the pre-trained monocular depth features to augment the final depth estimation, thereby improving the quality of Gaussian rendering.
@@ 3.1 Preliminary: Depth Foundation Model for Gaussian Splatting
For instance, DepthSplat processes multi-view images $\{I^{i}\}^{N}_{i=1}(I\in\mathbb{R}^{H\times W\times 3})$ using two parallel branches to extract dense per-pixel depth. One branch focuses on modeling cost volume features $C^{i}$ from the multi-view input, while the other employs a pre-trained monocular depth backbone, specifically Depth Anything V2, to obtain monocular depth features $F^{i}_{mono}$. Subsequently, the per-view cost volumes and monocular depth features are concatenated for 3D Gaussian prediction.
@@
3.2 Multi-modal Feature Matching
In this section, we seek to find an effective way to integrate sparse LiDAR depth into Gaussian Splatting with full utilization of multi-modal features. To this end, we propose a Multi-modal Feature Matching architecture tailored for urban scenarios to enable the synchronous integration of sparse depth information and color image data. Throughout this process, Depth-guided Position Embedding incorporates depth cues into the position embedding, enhancing 3D spatial awareness and improving multi-modal contextual comprehension.
Multi-modal feature matching.
As illustrated in Figure 2, at the core of our model is the Multi-modal Feature Matching of photometric features from the image and geometric cues from depth data. This is achieved through a Siamese-style encoder and an information cross-attention decoder, inspired by the DUSt3R series [30, 18].
Specifically, a monocular image $I\in\mathbb{R}^{H\times W\times 3}$ and a synchronized sparse depth map $S\in\mathbb{R}^{H\times W}$ are fed into a weight-sharing ViT encoder in a Siamese configuration, resulting in two token representations $F_{I}$ and $F_{S}$:
$F_{I}=\text{Encoder}(I),F_{S}=\text{Encoder}(S).$ (1)
$G_{I}$ $=\text{Decoder}_{1}(F_{I},F_{S}),$ (2)
$G_{S}$ $=\text{Decoder}_{2}(F_{S},F_{I}).$
Depth-guided positional embedding (DPE).
The conventional positional embedding in Vision Transformers encodes either relative or absolute spatial positions on a 2D image plane to ensure spatial awareness within the image. However, relying solely on the geometric properties of a 2D image plane is insufficient for our synchronized multi-modal design. To this end, we propose a straightforward depth-guided positional embedding (DPE) to integrate depth positions with image-based spatial positions.
Specifically, given the downsampled image size $H_{L}\times W_{L}$ and the sparse depth map, we first flatten the 2D grid of spatial positions into a 1D vector, where each element corresponds to a specific spatial location in the image. Subsequently, the sparse depth map is downsampled to match the image resolution, generating an independent set of depth indices that complement the spatial positions. The final positional embedding $D_{pos}$ is constructed by concatenating the flattened spatial positions with the depth positions, effectively encoding position information in the xy-z plane. By integrating both spatial and depth geometry, this module provides a comprehensive positional prior for effective multi-modal feature matching.
3.3 Multi-scale Gaussian Decoding
The two regression heads adhere to the DPT [22] architecture, enhanced with an additional multi-scale depth encoding that delivers precise scale priors for Gaussian prediction. In particular, at each scale within the DPT Decoder, we initially resize the input sparse depth map to align with the spatial size of the current feature scale. After that, the resized depth map is processed through a shallow network comprising two convolutional layers to extract depth features, which are then added to the DPT intermediate features. Finally, the input image and depth map, each processed by a single convolutional layer, are individually incorporated into the final features of the Gaussian Head and Geometry Head to facilitate either appearance-based or geometry-based Gaussian decoding.
3.4 Training Loss
Our model is trained using a combination of view synthesis loss and depth loss. The overall loss function is defined as follows:
$\mathcal{L}=\mathcal{L}_{\text{nvs}}+\mathcal{L}_{\text{depth}}$ (3)
Novel view synthesis loss.
where the LPIPS loss weight $\lambda$ is set to 0.05.
Depth loss.
4 Experiments
@@
Table 1: Quantitative comparison with state of the art on Waymo dataset. Our ADGaussian outperforms existing methods in nearly all scenarios. Cells highlighted in denotes the best and second-best performances.

### v2

# ADGaussian: Generalizable Gaussian Splatting for Autonomous Driving via Multi-modal Joint Learning

@@ Abstract
We present a novel approach, termed ADGaussian, for generalizable street scene reconstruction. The proposed method enables high-quality rendering from merely single-view input. Unlike prior Gaussian Splatting methods that primarily focus on geometry refinement, we emphasize the importance of joint optimization of image and depth features for accurate Gaussian prediction. To this end, we first incorporate sparse LiDAR depth as an additional input modality, formulating the Gaussian prediction process as a joint learning framework of visual information and geometric clue. Furthermore, we propose a Multi-modal Feature Matching strategy coupled with a Multi-scale Gaussian Decoding model to enhance the joint refinement of multi-modal features, thereby enabling efficient multi-modal Gaussian learning. Extensive experiments on Waymo and KITTI demonstrate that our ADGaussian achieves state-of-the-art performance and exhibits superior zero-shot generalization capabilities in novel-view shifting. Project page.
I INTRODUCTION
@@
Recently, 3D Gaussian Splatting (3DGS) [9] has garnered significant attention in the fields of 3D scene reconstruction and novel view synthesis [16] due to its real-time rendering speed and high-quality output. One key application is the modeling of street scenes from image sequences, which plays a vital role in areas such as autonomous driving [20].
Figure 1: We introduce ADGaussian, a generalizable Gaussian framework that achieves superior street scene reconstruction in both visual and geometric quality. The bottom row illustrates the results of viewpoint shifting, further demonstrating the robustness of our method under varying viewpoint changes.
@@
When modeling urban scenes, some methods follow per-scene optimization techniques [11, 33, 40], notably StreetGaussians [32] that represents dynamic urban street as a set of point clouds equipped with semantic logits and 3D Gaussians. While such an approach excels in high-quality reconstruction, it struggles with expensive training cost and large-range novel view synthesis, which motivates the exploration of generalizable models that avoid scene-specific fine-tuning.
@@
To achieve generalizable street scene reconstruction, most existing methods build upon the architectures of PixelSplat [1] or MVSplat [2]. For instance, GGRt [13] introduces a pose-free architecture to iteratively update multi-view depth map and subsequently estimates Gaussian primitives based on PixelSplat. Similarly, GGS [6] enhances the depth estimations of MVSplat by integrating a multi-view depth refinement module. Nevertheless, multi-view feature matching-based depth estimation may fail in challenging conditions such as texture-less areas and reflective surfaces. To tackle this issue, the concurrent work DepthSplat [31] combines pre-trained depth features from Depth Anything V2 [34] with multi-view depth estimations for accurate depth regression, where the estimated depth features are then concatenated with image features for Gaussian prediction.
@@
Given the great generalization capability of Depth Anything V2, it is reasonable to extend DepthSplat to urban street scenarios. However, DepthSplat faces specific limitations when applied to these environments. First, as highlighted in the dashed box of Fig. 1, the visual rendering quality is constrained by the performance of pre-trained depth models, which often exhibit inconsistent accuracy across diverse street datasets and scenarios. Second, simply concatenating image and depth features for the final Gaussian prediction without any information sharing or multi-modal feature fusion may lead to unexpected spatial misalignment, evidenced by the distorted car shape in the red box of Fig. 1.
@@
To overcome these limitations, we present ADGaussian, a novel multi-modal framework that jointly optimizes visual rendering quality and geometric accuracy for street scenes. Instead of relying on pre-trained depth foundation models, we choose to integrate sparse LiDAR depth information as an additional input modality, which is more practical in real-world scenarios and provides precise metric scale priors for geometry reconstruction. Given two complementary modalities (i.e., image data and sparse depth map), the framework’s core innovation lies in its synergistic processing of these two data streams, enabling effective information sharing and joint optimization between different modalities. Specifically, we introduce a Multi-modal Feature Matching strategy augmented by the Depth-guided Position Embedding, which contains a Siamese-style encoder paired with an information cross-attention decoder. This design ensures a cohesive fusion of geometric and appearance information, resulting in well-aligned multi-modal tokens. Subsequently, we employ a Multi-scale Gaussian Decoding model to aggregate multi-scale depth information into the resulting multi-modal tokens for the final 3D Gaussian predictions.
As presented in Fig. 1, ADGaussian excels in both visual rendering and geometric reconstruction. Notably, the bottom row demonstrates our model’s robustness under extreme viewpoint variations, where competing methods often produce distorted geometry. This superior performance validates the efficacy of ADGaussian’s synchronized cross-modal optimization paradigm.
@@
  • We present ADGaussian, the first generalizable framework that formulates street scene Gaussian prediction as joint visual and geometric learning.
@@
  • We conduct extensive comparisons on two datasets, verifying our approach’s state-of-the-art performance and the effectiveness of the proposed components.
II RELATED WORK
II-A Generalizable 3D Gaussian Splatting
@@
Generalizable Gaussian Splatting [28, 30, 3] aims to learn powerful priors that enable effective generalization across unseen scenes. Existing methods can be broadly categorized into two groups based on their handling of camera parameters. The first group, including approaches like MVSGaussian [14] and SplatterImage [22], predicts per-pixel 3D Gaussian primitives using known camera parameters. These works typically demonstrate superior reconstruction accuracy due to the precise geometric constraints provided by the camera poses. The second group of methods [19, 36] proposes to jointly predict camera parameters and 3D representations, eliminating the need for known camera poses. For instance, GGRt [13] employs an Iterative Pose Optimization Network to estimate and iteratively update the relative pose between target and reference images. DrivingForward [24] adopts pose network and depth network to determine the position of the Gaussian primitives in a self-supervised manner. In street scene modeling, however, camera poses provide critical constraints for determining scene scale and enhancing reconstruction accuracy from image sequences. Moreover, camera poses are readily accessible in street scenes, making them a practical and reliable data resource. Therefore, we choose to leverage posed images for our approach.
II-B Depth and Gaussian Splatting
@@
Depth quality has been demonstrated to play a pivotal role in Gaussian Splatting, serving as the foundation for accurate geometry reconstruction and realistic rendering. To ensure precise geometric fidelity, existing approaches [15, 41, 25] incorporate additional depth supervision into the optimization process. However, since dense Ground Truth depth data is often unavailable in practical applications, researchers have turned to pre-trained depth foundation models [17, 34] as an alternative source of reliable geometric cues. For example, Chung et al. [4] rescale pre-trained depth maps using sparse COLMAP points to provide depth constraints. DepthSplat [31] fuses pre-trained depth features with multi-view cost volume features to help depth refinement. While these methods have significantly improved geometric accuracy, we contend that their primary focus on geometry enhancement overlooks the crucial interplay between appearance and structure. In contrast to previous approaches, we argue that joint optimization of image and depth features is more critical for achieving high-quality reconstruction that excels in both visual fidelity and structural accuracy.
II-C LiDAR-Integrated Gaussian Splatting
@@
The integration of LiDAR data has emerged as a widely adopted approach in street scene reconstruction [26, 35], due to its effectiveness in facilitating geometry learning. The conventional methodologies typically involve two main steps: initializing Gaussians from LiDAR point clouds [10] to establish the basic scene structure, and further supervising predicted Gaussian positions with LiDAR priors [8, 7] to refine geometric details. Typically, TCLC-GS [39] constructs a hybrid 3D representation by combining LiDAR geometries with image colors, enabling simultaneous initialization of both geometric and appearance attributes of 3D Gaussians. Rather than directly using LiDAR point clouds, we propose leveraging sparse LiDAR depth to bridge the modality gap between LiDAR and camera data. Furthermore, we integrate depth priors as an additional input modality into a unified optimization framework, achieving joint optimization of depth geometries and image photometric attributes, as opposed to the common practice of initialization alone.
Figure 2: Overall framework of ADGaussian. Given monocular posed image and sparse depth as inputs, we first extract fused multi-modal features through Multi-modal Feature Matching strategy enhanced by Depth-guided Positional Embedding (DPE). These aligned multi-modal tokens are then processed by our Multi-scale Gaussian Decoding module, which hierarchically integrates depth cues across scales to produce optimized 3D Gaussian outputs.
III Method
@@
Depth foundation models [17, 34] have been integrated into the Gaussian Splatting to improve geometry reconstruction. However, such a framework often suffers from suboptimal rendering quality due to the insufficient interactions between photometric and geometric clues. To address this, we propose ADGaussian, a synchronized multi-modal optimization architecture that combines sparse depth data with monocular images for enhanced street scene modeling.
III-A Preliminary
@@
Recently, some works have investigated the advantages of using a pre-trained depth foundation model for image-conditioned 3D Gaussian reconstruction. For instance, DepthSplat [31] processes multi-view images $\{I^{i}\}^{N}_{i=1}(I\in\mathbb{R}^{H\times W\times 3})$ using two parallel branches to extract dense per-pixel depth. One branch focuses on modeling cost volume features $C^{i}$ from the multi-view input, while the other employs a pre-trained monocular depth backbone, specifically Depth Anything V2 [34], to obtain monocular depth features $F^{i}_{mono}$. These per-view cost volumes and monocular features are then concatenated for depth regression. Finally, DepthSplat predicts all remaining Gaussian parameters using the concatenated image, depth, and feature information.
@@
In this paper, instead of relying on pre-trained depth foundation models, we propose utilizing sparse LiDAR depth measurements as an additional input. This choice is particularly advantageous for street scene applications, as LiDAR data offers both greater practical accessibility in real-world autonomous driving systems and reliable metric-scale depth priors crucial for accurate geometric reconstruction.
III-B Multi-modal Feature Matching
In this subsection, we seek to find an effective way to integrate sparse LiDAR depth into Gaussian Splatting. To this end, we propose a Multi-modal Feature Matching architecture tailored for urban scenarios to enable the synchronous integration of sparse depth information and color image data. Throughout this process, Depth-guided Position Embedding incorporates depth cues into the position embedding, enhancing 3D spatial awareness and improving multi-modal contextual comprehension.
TABLE I: Quantitative comparisons with state of the art on Waymo dataset. Our ADGaussian outperforms existing methods in nearly all scenarios. Cells highlighted in denotes the best and second-best performances.
@@ III-B Multi-modal Feature Matching
Multi-modal Feature Matching
As illustrated in Fig. 2, the first part of our model is the Multi-modal Feature Matching of photometric features from the image and geometric cues from depth data. This is achieved through a Siamese-style encoder and an information cross-attention decoder, inspired by the DUSt3R series [27, 12].
Specifically, given a monocular image $I\in\mathbb{R}^{H\times W\times 3}$ and synchronized sparse depth map $S\in\mathbb{R}^{H\times W\times 1}$, we first replicate the depth map across channels to match the image’s dimensional structure. These multi-modal inputs are then fed into a weight-sharing ViT encoder, resulting in two token representations $F_{I}$ and $F_{S}$:
$F_{I}=\text{Encoder}(I),F_{S}=\text{Encoder}(S)$ (1)
$G_{I}$ $=\text{Decoder}_{1}(F_{I},F_{S}),$ (2)
$G_{S}$ $=\text{Decoder}_{2}(F_{S},F_{I})$
Depth-guided Positional Embedding (DPE)
The conventional positional embedding in Vision Transformers encodes either relative or absolute spatial positions on a 2D image plane to ensure spatial awareness within the image. However, relying solely on the geometric properties of a 2D image plane is insufficient for our synchronized multi-modal design. To this end, we propose a straightforward Depth-guided Positional Embedding (DPE) to integrate depth positions with image-based spatial positions
In particular, given the downsampled image size $H_{L}\times W_{L}$ and the sparse depth map, we first flatten the 2D grid of spatial positions into a 1D vector through row-major ordering, where each spatial position $(i,j)$ in the 2D grid is mapped to linear index $k=i\times W_{L}+j$ in the 1D vector. Meanwhile, the sparse depth map is downsampled to match the image resolution, generating an independent set of depth indices that complement the spatial positions. The final positional embedding $D_{pos}$ is then formed by concatenating the flattened spatial coordinates with their corresponding depth values, establishing an integrated xy-z coordinate representation that encodes both planar and depth-wise positional information. By integrating both spatial and depth geometry, this module provides a comprehensive positional prior for effective multi-modal feature fusion.
III-C Multi-scale Gaussian Decoding
The two regression heads adhere to the DPT [18] architecture, enhanced with an additional multi-scale depth encoding that delivers precise scale priors for Gaussian prediction. In particular, at each scale within the DPT Decoder, we initially resize the input depth map to align with the spatial size of the current feature scale. After that, the resized depth map is processed through a shallow network comprising two convolutional layers to extract depth features, which are then added to the DPT intermediate features. Finally, the input image and depth map, each processed by a single convolutional layer, are individually incorporated into the final features of the Gaussian Head and Geometry Head to facilitate either appearance-based or geometry-based Gaussian decoding.
III-D Training Loss
Our model is trained using a combination of novel view synthesis loss and depth loss:
$\mathcal{L}=\mathcal{L}_{\text{nvs}}+\mathcal{L}_{\text{depth}}$ (3)
Novel view synthesis loss
where the LPIPS loss weight $\lambda$ is set to 0.05.
Depth loss
IV Experiments
IV-A Implementation Details
Datasets and metrics
We evaluate our proposed approach on two widely used autonomous driving datasets: the Waymo Open Dataset [21] and the KITTI Tracking benchmarks [5]. For both datasets, we adopt a train-test split ratio of approximately 1:7. Specifically, on the Waymo dataset, our focus primarily lies on static and dynamic scenes, where each scene type is divided into 4 test scenes and 28 training scenes. Similarly, for the KITTI dataset, the split consists of 5 test scenes and 37 training scenes. For render quality evaluation, we employ the standard image quality metrics, including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [29], and the Learned Perceptual Image Patch Similarity (LPIPS) [37].
Figure 3: Qualitative comparisons with state of the art. Our ADGaussian surpasses all other competitive models in rendering quality within urban scenarios. The zoom-in comparisons in the first row further reveal our enhanced spatial alignment capabilities.
Training details
We employ the Adam optimizer and cosine learning rate schedule, with an initial learning rate of 1e1e-44. We train our model on a single 3090 Ti GPU, running for 150k iterations on both Waymo and KITTI datasets, with a batch size of 1. To ensure a fair comparison, all experiments are carried out at resolutions of 320×480 for the Waymo dataset and 256×608 for the KITTI dataset.
IV-B Comparisons with the State of the Art
When comparing our work with current state-of-the-art Gaussian Splatting methods, we selected MVSplat [2] (a multi-view cost volume-based approach) and DepthSplat [31] (a depth foundation model-based approach) as primary baselines. Since our method focuses on pose-aware generalizable Gaussian Splatting, we excluded pose-free methods and per-scene optimized baselines from the main comparisons. To ensure fair and consistent comparisons, all baseline methods were re-implemented and trained on both datasets using identical experimental settings to ours. Specifically, for each scenario, both MVSplat and DepthSplat utilize consecutive frame pairs as input, with the subsequent immediate frame serving as the target novel view for evaluation.
TABLE II: Quantitative comparisons with state of the art on KITTI dataset. KITTI¹ refers to subsequent temporal frame rendering, while KITTI² indicates spatial left-to-right camera rendering. Cells highlighted in denotes the best and second-best performances.
Method KITTI1 KITTI2